### v1

Abstract
We bring attention to the fact that the claim of Brambilla et.al. [1] of discovering a novel design for pentamode materials is incorrect. Back in 2016 Briane Harutyunyan and myself [2] designed a class of stiff pentamodes, that include the high bulk modulus pentamodes of Brambilla et.al.
Keywords: Stiff Pentamodes, High Bulk Modulus Pentamodes   Metal Water.
1
Pentamodes, as imaged in Figure 1, are a family of materials, introduced in [4] that have drawn considerable attention: see, for example, [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 3, 26, 27, 28, 29, 30].
Figure 1: An electron micrograph of the pentamode structure created by Kadic, Bückmann, Stenger, Thiel and Wegener [11] using a 3-dimensional lithography technique. (Used with the kind permission of Martin Wegener.)
Like liquids they support only one loading. Unlike liquids or gels, this loading need not be hydrostatic but rather could be a shear loading, or a combination of both hydrostatic and shear loadings. They are compliant to a set of 5 independent loadings (hence the name pentamode). Their elasticity tensor, when represented as a $6\times 6$ matrix, as is common in mechanical engineering, has 5 small eigenvalues and one much larger eigenvalue. The eigenvector corresponding to the large eigenvalue can be whatever one likes. Ideally, the ratios of the large eigenvalue to the small eigenvalues should be as large as possible. The original pentamode design had stiff double cone elements touching at points on a lattice and surrounded by a very compliant filler material. Alternatively, the double cone pivot junctions can be joined by narrow necks of the stiff phase and the filler phase can be void [11], thus making their construction more practical.
Pentamodes that are resistant to hydrostatic loadings include gels, pentamodes with double cones configured a diamond lattice arrangement [4], and the truss structure discovered independently by Sigmund [31] using topology optimization. A liquid is an ideal pentamode resistant to hydrostatic loading, but when placed on a flat surface it flows under gravity rather than keeping a desired shape.
For some applications, a disadvantage of the early pentamode designs with a given stiff phase is that as the pentamode design becomes more ideal (either by making the filler more and more compliant, or by narrowing the necks joining the double cone elements when they are surrounded by void) is that while the ratio of small to large eigenvalues of the elasticity tensor goes to zero, the (comparatively) large eigenvalue also goes to zero. One straightforward way to avoid this is to make the stiff phase increasingly stiff, in an appropriate manner, as the filler becomes more compliant or as the necks narrow. Brambilla et. al. [1] suggest that a more realistic alternative to obtain stiff pentamodes is to replace the double cones by cylindrical elements, one section of each contains an array of thin beams aligned with the cylinder axis. However, their claim that this construction is novel isn’t correct. The same class of stiff pentamodes was introduced by Briane, Harutyunyan and myself in 2016 [2]: see Figure 2.
Figure 2: Steps in the construction of stiff pentamodes as presented in [2], figures 9 and 10. In (a) the line segments between $P$ and the points $\bf{a_{i}},i=0,1,2,3$ are where the double cone structures are placed in the standard pentamode unit cell of periodicity. In (b) these line segments are replaced by cylinders of the stiff phase. (c) shows one such cylinder and two adjacent junctions. In (d) we replace the center section of each cylinder by a microstructure of parallel beams, as in (e), aligned with the cylinder axis.
The motivations were different. Brambilla et.al. suggested the design as a way to construct three dimensional metal-water, i.e., an elastically isotropic material with a stiff phase of metal surrounded by air having the bulk modulus of water and having an extremely small shear modulus. This is useful for acoustic cloaking [6, 32]. By contrast, our motivation was more theoretical: using them as a stepping stone towards partially characterizing the range of possible elasticity tensors achievable in composites of given proportions of a chosen isotropic stiff phase surrounded by an arbitrarily compliant filler phase. The stiff pentamodes were inserted as thin layers into materials that optimize sums of elastic energies under different loadings (the so-called Avellenada structures). These thin layers, with stiff pentamode substructures having unit cell size much smaller than the layer thickness, behave as slip planes undergoing infinitesimal slip in prescribed directions under infinitesimal elastic deformations. In particular, one can get a sequence of materials converging to the stiffest possible pentamode material built from the given phases, mixed in the desired volume fractions and supporting any single desired stress. These optimal pentamodes can be used to get a complete characterization of the set of all achievable (average stress, average strain) pairs in any composite of these two phases mixed in given proportions [3]. In two dimensions optimal pentamodes (technically bimode structures) supporting hydrostatic stresses were first discovered by Sigmund [33].
It is important to recognize that these constructions are only appropriate for linear elasticity: at any finite macroscopic deformation, the pentamode substructures in the thin layers will undergo large deformations and likely tear so that they no longer function as desired. The same is true, but to a lesser extent, of any of the pentamode structures. For example, under a finite deformation of pure shear the microstructure will stretch and cause the average stress to be non-zero. By contrast, there will be no such stress in a fluid as the deformation preserves volume.
Our routes for arriving at the beam substructure within cylindrical elements were distinctly different. Brambilla et.al. conceived it by analogy with a rope that is stiff when stretched or compressed, but which is easy to bend. By contrast, we realized that the beam substructure had the desired effective moduli: this was similar to my discovery in 1992 that the performance of a particular class of two-dimensional auxetic materials having a Poisson ratio approaching $-1$ could be improved by replacing the pivots with a beam (i.e., a two-dimensional laminate) substructure [34]: see Figure 3.
Figure 3: (a) A linkage used to keep the bottom and top surfaces parallel while allowing a sliding motion between them, as needed in the construction of a particular two-dimensional material having Poisson’s ratio close to -1, where the stiff phase is white and the highly compliant filler occupies the dotted regions. (b) to make it more realistic one can replace the pivot points by thin necks. (c) improved behavior in the limit of high contrast can be obtained by replacing the pivots with an array of beams (a two-dimensional laminate). This figure is adapted from figures 4 and 12 in [34].
By analogy with a rope, Brambilla et.al. also consider cylindrical elements with a substructure of beams that is then twisted, but they do not analyse this material. As they observe, such a structure has sliding surfaces and quite different responses under stretching and under compression. Therefore, neither their microscopic nor their macroscopic behaviors are that of linear continuum elasticity.
We emphasize that Brambilla et.al. have taken their analysis of the stiff pentamodes much further than us. Besides presenting beautiful renderings of the structure, they calculate the effective elasticity tensors when Ti 6Al-4V alloy is the stiff phase, they compute the band structure, and they identify the main low frequency modes of deformation.
Acknowledgments
The author is grateful to the National Science Foundation for support through Research Grant DMS-2107926.